Circular No. 8530

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2005ca AND 2005cb Further to IAUC 8513, M. SubbaRao, University of Chicago and Adler

Planetarium, on behalf of the SDSS Collaboration, reports the discovery of a supernova in spectra taken on May 12.19 UT. SN 2005ca is located at α = 13^{h}32^{m}38^{s}.59, δ = +11^{o}48′33′′.2 (equinox 2000.0), which is coincident with the center of the host galaxy whose redshift is z = 0.150. The new object appears to be a type-Ia supernova with an approximate age of 8 ± 5 days after maximum light, and its estimated magnitude is r = 20.9 ± 0.2 (the host galaxy has apparent mag r = 17.85 from an image taken prior to the supernova event on 2003 Mar. 31).

Further to IAUC 8499, C. Jacques reports the discovery of an apparent supernova (mag 15.6) by C. Colesanti, E. Pimentel, T. Napoleao and himself on unfiltered CCD images taken in the course of the Brazilian Supernovae Search on May 13.22 and 16.12 UT. SN 2005cb is located at α = 19^{h}11^{m}21^{s}.95, δ = −57^{o}02′27′′.4 (equinox 2000.0), which is 16′′ west and 19′′ north of the center of NGC 6753. Nothing in visible at this location on a CCD image taken on May 2.23 (limiting mag 18.5); a faint H II region is very close to the supernova position on that image and on a red Palomar Sky Survey image from 1976.

COMETS C/2005 F5 AND C/2005 G3–G6 (SOHO) Following are the initial available positions for additional Kreutz sun-

grazing comets found on SOHO website images (continuation to IAUC 8529; CL = C. Liang), with the following notes by K. Battams (as usual). C/2005 F5 was too diffuse to discern any tail; it reached mag ∼ 7.0 at 7.3R_{⊙} (Mar. 28.692). C/2005 G3 was stellar in appearance with no tail in the C3 images, reaching mag 6.2 at 9.4R_{⊙} (Apr. 4.279); C2 images showed a very short tail. C/2005 G4 was tiny and stellar in appearance, reaching roughly mag 7.5. C/2005 G5 and G6 showed no tail and were too diffuse for accurate photometry (though peak brightness was around mag 7).

Comet 2005 UT α_{2000} δ_{2000} Inst. F MPEC
C/2005 F5 Mar. 28.692 0^{h}37.^{m}9 + 2^{o}42′ C2  BZ 2005-J25
C/2005 G3 Apr. 3.904 1 07.5 + 3 29 C3/2 XL 2005-J25
C/2005 G4  8.160 1 15.4 + 6 20 C2  HS 2005-J25
C/2005 G5  9.935 1 22.8 + 7 09 C2  CL 2005-J25
C/2005 G6  11.685 1 29.7 + 7 44 C2  BZ 2005-J25
2005 May 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green